OpenStudy (anonymous):

Which test point holds true for 3/2y - 2x>1? A. (1/4, 1) B. (4, 1) C. (1, 1/4) D. (1, 0) E. (1/4, 3/4)

OpenStudy (anonymous):

i need help really bad

OpenStudy (anonymous):

lets try B

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

no no i am saying lets try it i have no idea if it is true or not

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

not true:(

OpenStudy (anonymous):

\[\frac{3}{2}y - 2x>1\] with \(x=4,y=1\) gives \[\frac{3}{2}\times 1-2\times 4>1\] or \(\frac{3}{2}-8>1\) which is wrong

OpenStudy (anonymous):

so it is b?

OpenStudy (anonymous):

no B is wrong

OpenStudy (anonymous):

lets try (1, 1/4)

OpenStudy (anonymous):

ok so which one should i go with?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have no idea you can't know it without trying them

OpenStudy (anonymous):

\[\frac{3}{2}\times \frac{1}{4}-2\times 1>1\] \[\frac{3}{8}-2>1\] no so it is not C

OpenStudy (anonymous):

ok so we have a. d. and e left

OpenStudy (anonymous):

\((1,0)\) does not work because \(0-2>1\) is false

OpenStudy (anonymous):

ok so i think a is right

OpenStudy (anonymous):

(1/4, 3/4)\[\frac{3}{2}\times \frac{3}{4}-2\times \frac{1}{4}>1\] \[\frac{9}{8}-\frac{1}{2}>1\] nope

OpenStudy (anonymous):

gosh so its e?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so its none of the above?

OpenStudy (anonymous):

maybe A (1/4, 1)\[\frac{3}{2}\times 1-2\times \frac{1}{4}>1\] \[\frac{3}{2}-\frac{1}{2}>1\] \[1>1\] hmm

OpenStudy (anonymous):

i don't see any that are right

OpenStudy (anonymous):

really:(?

OpenStudy (anonymous):

is it \(>\) or \(\ge\) ?

OpenStudy (anonymous):

lol then it is A